# Improvement and generalization of ABCD method with Bayesian inference

### v1

@@ 3.2.1 Small signal
To compare both methods, ABCD and Bayesian inference on a mixture model, we compare predicted and true signal events in each method. We do this in two different sets, one restricted to the signal region A, and the other in the full dataset. The first approach pursues to compare both methods within the rules of the ABCD method, in which the hypothesis is that the signal is localized in A. Since this assumption is usually not fully accomplished, we also study the second approach which is more relevant for observing signal events in general. We show both approaches for the three WP=70%, 80% and 90% in each plot in Fig. 7, where in both plots the horizontal and vertical axis are, respectively, the true and predicted number of signal events in the corresponding region.
@@
To interpret Fig. 7 with respect to the ABCD method, one should take into account what happens with signal and background events in the regions in Fig. 6 as the WP is increased. For a tight [^6] It is referred as tight because it practically does not tag non-b jets as b. WP (70%) we have that region A consists of almost all signal events with very few background events, however, at the price of leaving too many signal events in region C. Since the background in A is very small for this tight selection, any relative errors in background estimation due to contaminating signal events in C does not greatly affect the signal estimation in A. This is because the signal estimation is based on the total number of events in A minus background estimation in A and the former is much larger than the latter. Henceforth, the signal estimation for A with a tight WP is very good in the left plot in Fig. 7, but at the price of a very bad estimation of the total number of signal events in the sample, as depicted in the right plot in the same figure. Observe that one can easily understand the observed lower bias in the signal estimation in A because of the signal contamination in C.
@@
As one increases the WP, there is a migration of signal events from C to A which increases the localization of the signal, but also many more background events enter into all the regions since there are more $b$-tagged jets. The rate of population growth is different for background than for signal because the increase of false-positive $b$-tags from the non-b jets is larger than the increase in true-positive from the true b-jets. In particular, the background in A increases much more than the signal in A. Therefore, although there is a smaller relative bias in the background estimation in A –because of less contaminating signal events in C–, the increase of background relative to signal in A yields a more biased signal estimation in A. This can be seen in both plots in Fig. 7, and we have verified that the signal contamination in C is generating this bias. In any case, a looser WP (90%) yields a better estimation of the total number of signal events (right panel in figure) due to better localization.
@@
The notorious larger spread in the signal estimation as the WP increases is due to the same difference in growth between background and signal that biases the signal estimation. For tight WPs, the signal estimation is mostly the population in A corrected for a small background. Thus, the error in the signal estimation is mostly the usual Poisson error. The error in the ratio is thus approximately obtained by propagating the ratio between two Poisson variables, the signal estimation and the true signal. For larger WPs, because the signal-to-background fraction is smaller the signal estimation is more driven by the background estimation and its substraction from $N_{A}$. Thus, the error in the signal determination is more dominated by the errors in the background estimation. Because the errors in the background estimation are larger than the Poisson errors of each individual measurement, the uncertainty on the signal estimation increases noticeably more than the uncertainty on $N_{A}$ and on the true signal which also increase due to the increased WP. The uncertainty on the ratio between signal estimation and true signal increases accordingly.
@@
As it can be seen from the discussed figure, in all scenarios there is a better estimation from Bayesian inference than using the ABCD procedure. Actually, Bayesian inference does not rely on a selected WP, and therefore it is not affected by any problems associated with it. Of course, in the present work this is achievable thanks to the simplification in the parameterization of the $b$-tagging curve score and a more realistic case is discussed in Section 4. Also observe that the left plot in Fig. 7 has a dependence on the WP for Bayes solely because the signal region A has a dependence on it.
@@
Figure 7: Predicted versus true signal count in the ABCD signal region (left) and in the whole dataset (right). The observed biases and spreads in the ABCD predictions are discussed in the text. Dashed lines represent perfect matching, and we see that in all cases the Bayes framework compares favourably. See discussion in text concerning the ABCD performance when considering both left and right results. (Although Bayes does not depend on the WP, the ABCD signal region on which we are counting the events does depend on the WP. This is why there is a WP label assigned to the Bayesian method as well.)
@@ 3.2.2 No signal
Both for ABCD and the Bayes method, we obtain a usually non-integer signal prediction $S_{\mathrm{pred}}$. A more thorough analysis should assess the compatibility of this prediction with the actual presence of signal events, which can only take discrete values. This could be done by using a Poisson distribution with rate given by the estimated $S_{\mathrm{pred}}$ and evaluating the probability of zero events. Another possibility is to study the compatibility of the data with the $\pi_{s}=0$ hypothesis. This could be done in a fully Bayesian manner by computing the evidence ratio [13] between the two probabilistic models (with and without signal) or approximately by re-doing the MAP estimation with two classes and computing a model comparison metric such as the Bayesian Information Criterion difference between models [13]. However, the presented analysis already provides a satisfactory proof of robustness and we leave these more involved model comparison techniques for future work.

### v2

@@ 3.2.1 Small signal
To compare both methods, ABCD and Bayesian inference on a mixture model, we compare predicted and true signal events in each method. We do this in two different sets, one restricted to the signal region A, and the other in the full dataset. The first approach pursues to compare both methods within the rules of the ABCD method, in which the hypothesis is that the signal is localized in A. Since this assumption is usually not fully accomplished, we also study the second approach which is more relevant for observing signal events in general and can highlight the known ABCD pitfalls when small signal fluctuations bias the background estimation. We show both approaches for the three WP=70%, 80% and 90% in each plot in Fig. 7, where we show the predicted signal events and the difference between predicted and true signal events as a function the true number of signal events in the corresponding region.
@@
To interpret Fig. 7 with respect to the ABCD method, one should take into account how the signal and background events are redistributed among the regions in Fig. 6 as the WP is increased. For a tight [^6] It is referred as tight because it practically does not tag non-b jets as b. WP (70%) we have that region A consists of almost all signal events with very few background events, however, at the price of leaving too many signal events in region C. Since the background in A is very small for this tight selection, any relative errors in background estimation due to contaminating signal events in C does not greatly affect the signal estimation in A. This is because the signal estimation is based on the total number of events in A minus background estimation in A and the former is much larger than the latter. Henceforth, the signal estimation for A with a tight WP is very good in the left column in Fig. 7, but at the price of a very bad estimation of the total number of signal events in the sample, as depicted in the right column in the same figure. Observe that one can easily understand the observed lower bias in the signal estimation in A because of the signal contamination in C.
@@
As one increases the WP, there is a migration of signal events from C to A which increases the localization of the signal, but also many more background events enter into all the regions since there are more $b$-tagged jets. The rate of population growth is different for background than for signal because the increase of false-positive $b$-tags from the non-b jets is larger than the increase in true-positive from the true b-jets. In particular, the background in A increases noticeably more than the signal in A. Therefore, although there is a smaller relative bias in the background estimation in A –because of less contaminating signal events in C–, the increase of background relative to signal in A yields a more biased signal estimation in A. This can be seen in both columns in Fig. 7, and we have verified that the signal contamination in C is generating this bias. In any case, a looser WP (90%) yields a better estimation of the total number of signal events (right column in figure) due to better localization.
@@
The notorious larger spread in the signal estimation as the WP increases is due to the same difference in growth between background and signal that biases the signal estimation. For tight WPs, the signal estimation is mostly the population in A corrected for a small background. Thus, the error in the signal estimation is mostly the usual Poisson error. The error in the ratio is thus approximately obtained by propagating the ratio between two Poisson variables, the signal estimation and the true signal. For larger WPs, because the signal-to-background fraction is smaller the signal estimation is driven by the background estimation and its subtraction from $N_{A}$. Thus, the error in the signal determination is more dominated by the errors in the background estimation. Because the errors in the background estimation are larger than the Poisson errors of each individual measurement, the uncertainty on the signal estimation increases noticeably more than the uncertainty on $N_{A}$ and on the true signal which also increase due to the increased WP. The uncertainty on the ratio between signal estimation and true signal increases accordingly.
@@
As it can be seen from the discussed figure, in all scenarios there is a better estimation from Bayesian inference than using the ABCD procedure. Because Bayesian inference does not rely on a selected WP it is therefore not affected by any problems associated with it. Of course, in the present work this is achievable thanks to the simplification in the parameterization of the $b$-tagging curve score and a more realistic case is discussed in Section 4. Also observe that the left plot in Fig. 7 has a dependence on the WP for Bayes solely because the signal region A has a dependence on it.
@@
Figure 7: Top (bottom): Predicted signal count (difference between predicted and true signal counts) as a function of true signal count in the ABCD signal region (left) and in the whole dataset (right). The observed biases and spreads in the ABCD predictions are discussed in the text. Dashed lines represent perfect matching, and we see that in all cases the Bayes framework compares favourably. See discussion in text concerning the ABCD performance when considering both left and right results. (Although Bayes does not depend on the WP, the ABCD signal region on which we are counting the events does depend on the WP. This is why there is a WP label assigned to the Bayesian method as well.)
@@ 3.2.2 No signal
Both for ABCD and the Bayes method, we obtain a usually non-integer signal prediction $S_{\mathrm{pred}}$. A more thorough analysis should assess the compatibility of this prediction with the actual presence of signal events, which can only take discrete values. This could be done by using a Poisson distribution with rate given by the estimated $S_{\mathrm{pred}}$ and evaluating the probability of zero events. Another possibility is to study the compatibility of the data with the $\pi_{s}=0$ hypothesis. This could be done in a fully Bayesian manner by computing the evidence ratio (see Section 3.4 in  [15]) between the two probabilistic models (with and without signal), performing a posterior predictive check [22] for each model; or approximately by re-doing the MAP estimation with two classes and computing a model comparison metric such as the Bayesian Information Criterion difference between models (see Section 3.5 in  [15]). However, the presented analysis already provides a satisfactory proof of robustness and we leave these more involved model comparison techniques for future work.